OpenStudy (anonymous):

Given the function f(x)=(x-12)^2 where x less than or equal to 12 , find f-1(x) and its domain.

OpenStudy (anonymous):

sqrt(y) = x - 12 x = sqrt(y) + 12

OpenStudy (anonymous):

so inverse is y = sqrt(x) + 12

OpenStudy (anonymous):

it's domain will be x >= 0

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

wait so does that mean [0,inf)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

domain is right. web work keeps saying the answer, the inverse function that is, is wrong

OpenStudy (tkhunny):

It's always good to find the Domain and Range of \(f(x)\) before you start. This makes the Range and Domain of \(f^{-1}(x)\) rather trivial. Anyway, let's try that swap-and-solve again. If \(y = (x-12)^{2}\) We have \(x = (y-12)^{2}\) to solve for \(y\) The square root - be very carful, here \(\pm\sqrt{x} = y-12\) Finally, \(y = 12 \pm \sqrt{x}\) Notice how there are TWO answers. The challenge of this problem is to pick the correct branch. Hint: Whe had the Domain and Range of \(f(x)\). Which branch reflects those original properties?